OpenStudy (shaleiah):

g(3) g(x)=x^2-7x+1= Help me evaluate? (step by step)

OpenStudy (khantahmina):

plug 3 in where there is an x

OpenStudy (shaleiah):

I get -20

OpenStudy (fluttershyk):

3x- x^2+7x-1=0 10x- x^2-1=0 x= -10+4√6/-2, -10- 4 √6/-2 x=5-2 √6, 5+2 √6

OpenStudy (shaleiah):

@KendrickLamar2014

OpenStudy (khantahmina):

i got -11

OpenStudy (shaleiah):

could you guide me step by step?

OpenStudy (khantahmina):

open this

OpenStudy (shaleiah):

The instructor wants me to do it step by step.

OpenStudy (shaleiah):

When I plug in 3, I get -20

OpenStudy (khantahmina):